# Brain network science modelling of sparse neural networks enables Transformers and LLMs to perform as fully connected

### v2

@@ Abstract
This study aims to enlarge our current knowledge on the application of brain-inspired network science principles for training artificial neural networks (ANNs) with sparse connectivity. Dynamic sparse training (DST) emulates the synaptic turnover of real brain networks, reducing the computational demands of training and inference in ANNs. However, existing DST methods face difficulties in maintaining peak performance at high connectivity sparsity levels. The Cannistraci-Hebb training (CHT) is a brain-inspired method that is used in DST for growing synaptic connectivity in sparse neural networks. CHT leverages a gradient-free, topology-driven link regrowth mechanism, which has been shown to achieve ultra-sparse (1% connectivity or lower) advantage across various tasks compared to fully connected networks. Yet, CHT suffers two main drawbacks: (i) its time complexity is $\mathcal{O}(N\cdot d^{3})$- N node network size, d node degree - hence it can be efficiently applied only to ultra-sparse networks. (ii) it rigidly selects top link prediction scores, which is inappropriate for the early training epochs, when the network topology presents many unreliable connections. Here, we design the first brain-inspired network model - termed bipartite receptive field (BRF) - to initialize the connectivity of sparse artificial neural networks. Then, we propose a matrix multiplication GPU-friendly approximation of the CH link predictor, which reduces the computational complexity to $\mathcal{O}(N^{3})$, enabling a fast implementation of link prediction in large-scale models. Moreover, we introduce the Cannistraci-Hebb training soft rule (CHTs), which adopts a flexible strategy for sampling connections in both link removal and regrowth, balancing the exploration and exploitation of network topology. Additionally, we propose a sigmoid-based gradual density decay strategy, leading to an advanced framework referred to as CHTss. Empirical results show that BRF offers performance advantages over previous network science models. Using 1% of connections, CHTs outperforms fully connected networks in MLP architectures on visual classification tasks, compressing some networks to less than 30% of the nodes. Using 5% of the connections, CHTss outperforms fully connected networks in two Transformer-based machine translation tasks. Finally, using 30% of the connections, CHTs and CHTss achieve superior performance compared to other dynamic sparse training methods in language modeling across different sparsity levels, and it surpasses the fully connected counterpart in zero-shot evaluations.
@@ 1 Introduction
To evaluate the effectiveness of the Cannistraci-Hebb Training soft rule with sigmoid density decay (CHTss), we conduct extensive experiments across multiple architectures and tasks. Firstly, to assess the basic concept of CHTs, we employ MLPs on benchmark datasets, including MNIST MNIST , EMNIST cohen2017emnist , and Fashion MNIST xiao2017fashion . The results show that CHTs performs better than fully connected networks with only 1% of the connections (99% sparsity) in MLPs. Further, to evaluate the end-to-end approach CHTss, we utilize Transformers Transformer on machine translation datasets such as Multi30k en-de multi30k , IWSLT14 en-de cettolo-etal-2014-report , and WMT17 en-de bojar2017wmt . Additionally, we test CHTs and CHTss also on the LLaMA-130M model with language modeling tasks and zero-shot evaluation tasks. From the results of the above experiments, CHTss outperforms fully connected Transformers with only 5% of the links on Multi30k and IWSLT and achieves performance comparable to the fully connected LLaMA-130M in language modeling tasks on OpenWebText. Moreover, CHTs and CHTss outperform the fully connected LLaMA-130M on zero-shot evaluation tasks on GLUE wang2019glue and SuperGLUE wang2020supergluestickierbenchmarkgeneralpurpose with only 30% connections. These findings underscore the potential of CHTs and CHTss in enabling highly efficient and effective large-scale sparse neural network training.
@@ 2.1 Dynamic sparse training
Dynamic sparse training is a subset of sparse training methodologies. Unlike static sparse training methods (also known as pruning at initialization) prabhu2018deep ; SNIP ; dao2022monarch ; stewart2023data , dynamic sparse training allows for the evolution of network topology during the training process. The pioneering method in this field is Sparse Evolutionary Training (SET) SET , which removes links based on the magnitude of their weights and regrows new links randomly. Subsequent developments have sought to refine and expand upon this concept of dynamic topological evolution. One such advancement was proposed by DeepR bellec2017deep , a method that adjusts network connections based on stochastic gradient updates combined with a Bayesian-inspired update rule. Another significant contribution is RigL RIGL , which leverages the gradient information of non-existing links to guide the regrowth of new connections during training. MEST MEST utilizes both gradient and weight magnitude information to selectively remove and randomly regrow new links, analogously to SET. In addition, it introduces an EM&S strategy that allows the model to train at a higher density and gradually converge to the target sparsity. The Top-KAST jayakumar2020top method maintains constant sparsity throughout training by selecting the top $K$ parameters based on parameter magnitude at each training step and applying gradients to a broader subset $B$, where $B\supset A$. To avoid settling on a suboptimal sparse subset, Top-KAST also introduces an auxiliary exploration loss that encourages ongoing adaptation of the mask. Additionally, sRigL lasby2023srigl adapts the principles of RigL to semi-structured sparsity, facilitating the training of vision models from scratch with actual speed-ups during training phases. Despite these advancements, the state-of-the-art method remains RigL-based, yet it is not fully sparse in backpropagation, necessitating the computation of gradients for non-existing links. Addressing this limitation, Zhang et al. zhang2023ultra propose CHT, a dynamic sparse training methodology that adopts a gradient-free regrowth strategy that relies solely on topological information (network shape intelligence), achieving an ultra-sparse configuration that surpasses fully connected networks in some tasks.
@@ 2.2 Cannistraci-Hebb Theory and Network Shape Intelligence
As the SOTA gradient-free link regrown method, CHT zhang2023ultra originates from a brain-inspired network science theory. Drawn from neurobiology, Hebbian learning was introduced in 1949 (hebbian, ) and can be summarized in the axiom: “neurons that fire together wire together.” This could be interpreted in two ways: changing the synaptic weights (weight plasticity) and changing the shape of synaptic connectivity (cannistraci2013link, ; 18, ; 19, ; 20, ; Narula, ). The latter is also called epitopological plasticity (cannistraci2013link, ) because plasticity means “to change shape,” and epitopological means “via a new topology.” Epitopological Learning (EL) (18, ; 19, ; 20, ) is derived from this second interpretation of Hebbian learning and studies how to implement learning on networks by changing the shape of their connectivity structure. One way to implement EL is via link prediction, which predicts the existence and likelihood of each nonobserved link in a network. CH3-L3 is one of the best-performing and most robust network automata, belonging to the Cannistraci-Hebb (CH) theory (CHA, ), which can automatically evolve the network topology starting from a given structure. The rationale is that, in any complex network with local-community organization, the cohort of nodes tends to be co-activated (fire together) and to learn by forming new connections between them (wire together) because they are topologically isolated in the same local community (CHA, ). This minimization of the external links induces a topological isolation of the local community, which is equivalent to forming a barrier around it. The external barrier is fundamental to maintaining and reinforcing the signaling in the local community, inducing the formation of new links that participate in epitopological learning and plasticity.
@@ 3 Cannistraci-Hebb training soft rule with sigmoid gradual density decay
In this article, we adopt a probabilistic approach where the regrowth process is modeled as sampling from a $\{0,1\}$ multinomial distribution, with probabilities determined by link prediction scores, thereby introducing a "soft sampling" mechanism. Under this formulation, each mask value is not rigidly dictated by the link prediction score; instead, low-score links can still be selected with lower probability, facilitating escape from epitopological local minima (ELM). To demonstrate that soft sampling effectively balances exploration and exploitation, we evaluate its behavior in Figure 3, which presents the impact of varying softness levels during training of LLaMA-60M for 5000 steps under 90% sparsity.
We compare the in-time over-parameterization (ITOP) rate liu2021itop , which quantifies the cumulative proportion of links activated throughout training. As shown, deterministic regrowth leads to rapid topological convergence after approximately 1000 steps, indicating that it becomes trapped in an ELM without further exploration. Random regrowth, while capable of escaping ELMs by introducing new connections, fails to exploit the underlying topological structure effectively. In contrast, soft regrowth achieves a balance by both exploiting the current topology and exploring new link combinations probabilistically. This balance enables a more appropriate exploration of the connectivity space, ultimately leading to superior performance, as evidenced by the results.
@@
[…]
@@
3.3 Sigmoid Gradual Decrease Density
@@
However, this density decay scheduler exhibits a sharp decline in the initial stages of training, which risks pruning a substantial fraction of weights before the model has sufficiently learned. To mitigate this issue, we propose a sigmoid-based gradual density decrease strategy, defined as:
@@
where $t\in\{t_{0},t_{0}+\Delta t,\ldots,t_{0}+n\Delta t\}$, $s_{i}$ is the initial sparsity, $s_{f}$ is the target sparsity, $t_{0}$ is the starting epoch of gradual pruning, $t_{f}$ is the end epoch of gradual pruning, and $\Delta t$ is the pruning frequency. $k$ controls the curvature of the decrease. We set $k$=6 for all the experiments in this article. This strategy ensures a smoother initial pruning phase, allowing the model to warm up and stabilize before undergoing significant pruning, thereby enhancing training stability and performance. We explain how to adjust the training FLOPs of sigmoid density decay to the same as cubic decay in Appendix I.
@@ 4 Experiments
Table 1: Performance comparison of different fixed sparsity dynamic sparse training methods on the CIFAR10 dataset trained on an MLP at 99% sparsity. ACC represents accuracy, and ANP denotes the active neuron percolation rate, indicating the final size of the network. The lowest anp rate and the best dynamic sparse training method are highlighted in bold, and performances surpassing the fully connected model are marked with “*". The results present a standard error taken over three seeds of the experiments.
@@ 4.1 Experimental Setup
We evaluate the performance of CHTs using MLPs for image classification tasks on the MNIST MNIST , Fashion MNIST xiao2017fashion , EMNIST cohen2017emnist , and CIFAR10 CIFAR10 datasets. To further validate our approach, we apply the sigmoid gradual density decay strategy to Transformers for machine translation tasks on the Multi30k en-de multi30k , IWSLT14 en-de cettolo-etal-2014-report , and WMT17 en-de bojar2017wmt datasets. Additionally, we conduct language modeling experiments using the OpenWebText dataset Gokaslan2019OpenWeb and evaluate zero-shot performance on the GLUE wang2019glue and SuperGLUE wang2020supergluestickierbenchmarkgeneralpurpose benchmark with LLaMA-130M touvron2023llama . For MLP training, we sparsify all layers except the final layer, as ultra-sparsity in the output layer may lead to disconnected neurons, and the connections in the final layer are relatively minor compared to the previous layers. For Transformers and LLaMA models, we apply dynamic sparse training (DST) to all linear layers, excluding the embedding and final generator layer. Detailed hyperparameter settings for each experiment are provided in Tables 56, and 7. We also conduct a series of ablation and sensitivity tests on the components proposed in this article for CHTs and CHTss in Appendix H
@@ 4.2 MLP for image classification
Table 2: Performance comparison on machine translation tasks of Multi30k, IWSLT, and WMT with varying final sparsity levels. The scores indicate BLEU scores, which is the higher the better. CHTs (GMP) represents CHTs with GMP’s density decay strategy. Bold values denote the best performance among fixed sparsity DST methods or density decay DST methods. The performances that surpass the fully connected model are marked with “*". $s_{i}$ indicates the starting sparsity of the DST methods that use the density decay strategy.
@@
CHTs 29.45 29.89 21.39 22.02 21.05 22.36
MESTEM&S 28.71 28.26 18.95 20.77 20.79 22.3
GMP ($s_{i}=0.5$) 26.42 27.06 22.44 22.62 22.29 23.52
GraNet ($s_{i}=0.5$) 30.90 31.06 23.05 22.88 22.11 23.49
CHTss ($s_{i}=0.5$) 32.82* 33.11* 24.84* 24.76* 22.84 24.80
@@
In the MLP evaluation, we aim to assess the fundamental capacity of DST methods to train the fully connected module, which is common across many ANNs. The sparse topological initialization of CHT and CHTs is CSTI zhang2023ultra since the input bipartite layer can directly receive information from the input pixels. Table 8 displays the performance of DST methods compared to their fully connected counterparts across three basic datasets of MNIST, Fashion MNIST, and EMNIST. The DST methods are tested at 99% sparsity. As shown in Table 8, both of the two regrowth methods of CHTs outperform the other fixed sparsity DST methods. Notably, the path-based CH3-L3p outperforms the fully connected one in all the datasets. The node-based CH2-L3n also achieves comparable performance on these basic datasets. However, considering the running time of CH3-L3p is unacceptable, especially in large scale models, in the rest of the experiments of this article, we only use CH2-L3n as the representative method to regrow new links. Table 1 presents a comparison of fixed-sparsity dynamic sparse training (DST) methods against the fully connected (FC) baseline. Notably, CHTs outperform all other DST methods and achieve an 11% improvement in accuracy over the fully connected model. In addition, we present the active neuron post-percolation rate (ANP) for each method in Table 8 and Table 1. It is evident that CHTs adaptively percolates the network more effectively while retaining performance.

### v3

@@ Abstract
This study aims to enlarge our current knowledge on the application of brain-inspired network science principles for training artificial neural networks (ANNs) with sparse connectivity. Dynamic sparse training (DST) emulates the synaptic turnover of real brain networks, reducing the computational demands of training and inference in ANNs. However, existing DST methods face difficulties in maintaining peak performance at high connectivity sparsity levels. The Cannistraci-Hebb training (CHT) is a brain-inspired method that is used in DST for growing synaptic connectivity in sparse neural networks. CHT leverages a gradient-free, topology-driven link regrowth mechanism, which has been shown to achieve ultra-sparse (1% connectivity or lower) advantage across various tasks compared to fully connected networks. Yet, CHT suffers two main drawbacks: (i) its time complexity is $\mathcal{O}(N\cdot d^{3})$- N node network size, d node degree - hence it can be efficiently applied only to ultra-sparse networks. (ii) it rigidly selects top link prediction scores, which is inappropriate for the early training epochs, when the network topology presents many unreliable connections. Here, we design the first brain-inspired network model - termed bipartite receptive field (BRF) - to initialize the connectivity of sparse artificial neural networks. Then, we propose a matrix multiplication GPU-friendly approximation of the CH link predictor, which reduces the computational complexity to $\mathcal{O}(N^{3})$, enabling a fast implementation of link prediction in large-scale models. Moreover, we introduce the Cannistraci-Hebb training soft rule (CHTs), which adopts a flexible strategy for sampling connections in both link removal and regrowth, balancing the exploration and exploitation of network topology. Additionally, we propose a sigmoid-based gradual density decay strategy, leading to an advanced framework referred to as CHTss. Empirical results show that BRF offers performance advantages over previous network science models. Using 1% of connections, CHTs outperforms fully connected networks in MLP architectures on visual classification tasks, compressing some networks to less than 30% of the nodes. Using 5% of the connections, CHTss outperforms fully connected networks in two Transformer-based machine translation tasks. Finally, with only 30% of the connections, both CHTs and CHTss achieve superior performance over other dynamic sparse training methods, and perform on par with—or even surpass—their fully connected counterparts in language modeling across various sparsity levels within the LLaMA model family. The code is available at: https://github.com/biomedical-cybernetics/Cannistraci-Hebb-training.
@@ 1 Introduction
To evaluate the effectiveness of the Cannistraci-Hebb Training soft rule with sigmoid density decay (CHTss), we conduct extensive experiments across multiple architectures and tasks. Firstly, to assess the basic concept of CHTs, we employ MLPs on benchmark datasets, including MNIST MNIST, EMNIST cohen2017emnist, and Fashion MNIST xiao2017fashion. The results show that CHTs performs better than fully connected networks with only 1% of the connections (99% sparsity) in MLPs. Further, to evaluate the end-to-end approach CHTss, we utilize Transformers Transformer on machine translation datasets such as Multi30k en-de multi30k, IWSLT14 en-de cettolo-etal-2014-report, and WMT17 en-de bojar2017wmt. From the experimental results, CHTss surpasses fully connected Transformers while using only 5% of the connections on Multi30k and IWSLT, and achieves performance comparable to fully connected LLaMA-60M and LLaMA-130M models on OpenWebText language modeling tasks. Moreover, with 30% of the connections, CHTs even outperforms the fully connected LLaMA-1B counterpart. These findings underscore the potential of CHTs and CHTss in enabling highly efficient and effective large-scale sparse neural network training.
@@ 2.1 Dynamic sparse training
Dynamic sparse training is a subset of sparse training methodologies. Unlike static sparse training methods (also known as pruning at initialization) prabhu2018deep; SNIP; dao2022monarch; stewart2023data, dynamic sparse training allows for the evolution of network topology during the training process. The pioneering method in this field is Sparse Evolutionary Training (SET) SET, which removes links based on the magnitude of their weights and regrows new links randomly. Subsequent developments have sought to refine and expand upon this concept of dynamic topological evolution. One such advancement was proposed by DeepR bellec2017deep, a method that adjusts network connections based on stochastic gradient updates combined with a Bayesian-inspired update rule. Another significant contribution is RigL RIGL, which leverages the gradient information of non-existing links to guide the regrowth of new connections during training. MEST MEST utilizes both gradient and weight magnitude information to selectively remove and randomly regrow new links, analogously to SET. In addition, it introduces an EM&S strategy that allows the model to train at a higher density and gradually converge to the target sparsity. The Top-KAST jayakumar2020top method maintains constant sparsity throughout training by selecting the top $K$ parameters based on parameter magnitude at each training step and applying gradients to a broader subset $B$, where $B\supset A$. To avoid settling on a suboptimal sparse subset, Top-KAST also introduces an auxiliary exploration loss that encourages ongoing adaptation of the mask. Additionally, sRigL lasby2023srigl adapts the principles of RigL to semi-structured sparsity, facilitating the training of vision models from scratch with actual speed-ups during training phases. Despite these advancements, the state-of-the-art method remains RigL-based, yet it is not fully sparse in backpropagation, necessitating the computation of gradients for non-existing links. Addressing this limitation, Zhang et al. zhang2023ultra propose CHT, a dynamic sparse training methodology that adopts a gradient-free regrowth strategy that relies solely on topological information (network shape intelligence), achieving an ultra-sparse configuration that surpasses fully connected networks in some tasks. Some extra related works of DST are provided in Appendix M.
@@ 2.2 Cannistraci-Hebb Theory and Network Shape Intelligence
As the SOTA gradient-free link regrown method, CHT zhang2023ultra originates from a brain-inspired network science theory. Drawn from neurobiology, Hebbian learning was introduced in 1949 (hebbian) and can be summarized in the axiom: “neurons that fire together wire together.” This could be interpreted in two ways: changing the synaptic weights (weight plasticity) and changing the shape of synaptic connectivity (cannistraci2013link; 18; 19; 20; Narula). The latter is also called epitopological plasticity (cannistraci2013link) because plasticity means “to change shape,” and epitopological means “via a new topology.” Epitopological Learning (EL) (18; 19; 20) is derived from this second interpretation of Hebbian learning and studies how to implement learning on networks by changing the shape of their connectivity structure. One way to implement EL is via link prediction, which predicts the existence and likelihood of each nonobserved link in a network. CH3-L3 is one of the best-performing and most robust network automata, belonging to the Cannistraci-Hebb (CH) theory (CHA), which can automatically evolve the network topology starting from a given structure. The rationale is that, in any complex network with local-community organization, the cohort of nodes tends to be co-activated (fire together) and to learn by forming new connections between them (wire together) because they are topologically isolated in the same local community (CHA). This minimization of the external links induces a topological isolation of the local community, which is equivalent to forming a barrier around it. The external barrier is fundamental to maintaining and reinforcing the signaling in the local community, inducing the formation of new links that participate in epitopological learning and plasticity.
@@ 3 Cannistraci-Hebb training soft rule with sigmoid gradual density decay
In this article, we adopt a probabilistic approach where the regrowth process is modeled as sampling from a $\{0,1\}$ multinomial distribution, with probabilities determined by link prediction scores, thereby introducing a "soft sampling" mechanism. Under this formulation, each mask value is not rigidly dictated by the link prediction score; instead, low-score links can still be selected with lower probability, facilitating escape from epitopological local minima (ELM).
To demonstrate that soft sampling effectively balances exploration and exploitation, we evaluate its behavior in Figure 3, which presents the impact of varying softness levels during training of LLaMA-60M for 5000 steps under 90% sparsity. We compare the in-time over-parameterization (ITOP) rate liu2021itop, which quantifies the cumulative proportion of links activated throughout training. As shown, deterministic regrowth leads to rapid topological convergence after approximately 1000 steps, indicating that it becomes trapped in an ELM without further exploration. Random regrowth, while capable of escaping ELMs by introducing new connections, fails to exploit the underlying topological structure effectively. In contrast, soft regrowth achieves a balance by both exploiting the current topology and exploring new link combinations probabilistically. This balance enables a more appropriate exploration of the connectivity space, ultimately leading to superior performance, as evidenced by the results.
@@
We illustrate the link removal part of CHTs in Figure 1b1) and b2). We employ two methods, Weight Magnitude (WM) $|\textbf{W}|$ and Relative Importance (RI) zhangplug, to remove the connections during dynamic sparse training. Given an input node $i$ and an output node $j$ connected with weight $W_{ij}$, we define the relative importance $RI_{ij}$ as:
@@
3.3 Sigmoid Gradual Density Decay
@@
However, this density decay scheduler exhibits a sharp decline in the initial stages of training, which risks pruning a substantial fraction of weights before the model has sufficiently learned. To enable a smoother decay during the pruning stage, we propose a sigmoid-based gradual density decrease strategy, defined as:
@@
where $t\in\{t_{0},t_{0}+\Delta t,\ldots,t_{0}+n\Delta t\}$, $s_{i}$ is the initial sparsity, $s_{f}$ is the target sparsity, $t_{0}$ is the starting epoch of gradual pruning, $t_{f}$ is the end epoch of gradual pruning, and $\Delta t$ is the pruning frequency. $k$ controls the curvature of the decrease. We set kk=6 for all the experiments in this article. This strategy ensures a smoother initial pruning phase, allowing the model to warm up and stabilize before undergoing significant pruning, thereby enhancing training stability and performance. We explain how to adjust the training FLOPs of sigmoid density decay to the same as cubic decay in Appendix I.
@@ 4 Experiments
Table 1: Performance comparison of different sparsity dynamic sparse training methods on the CIFAR10 dataset trained on an MLP at 99% sparsity. The density decay of GMP, GraNet, and CHTss starts with a sparsity of 50%. ACC represents accuracy, and ANP denotes the active neuron percolation rate, indicating the final size of the network. The lowest anp rate and the best dynamic sparse training method are highlighted in bold, and performances surpassing the fully connected model are marked with “*". The results present a standard error taken over three seeds of the experiments.
@@ 4.1 Experimental Setup
We evaluate the performance of CHTs using MLPs for image classification tasks on the MNIST MNIST, Fashion MNIST xiao2017fashion, EMNIST cohen2017emnist, and CIFAR10 CIFAR10 datasets. To further validate our approach, we apply the sigmoid gradual density decay strategy to Transformers for machine translation tasks on the Multi30k en-de multi30k, IWSLT14 en-de cettolo-etal-2014-report, and WMT17 en-de bojar2017wmt datasets. Additionally, we conduct language modeling experiments using the OpenWebText dataset Gokaslan2019OpenWeb and evaluate zero-shot performance on the GLUE wang2019glue and SuperGLUE wang2020supergluestickierbenchmarkgeneralpurpose benchmark with LLaMA-60M, 130M, and 1B models touvron2023llama. For MLP training, we sparsify all layers except the final layer, as ultra-sparsity in the output layer may lead to disconnected neurons, and the connections in the final layer are relatively minor compared to the previous layers. For Transformers and LLaMA models, we apply dynamic sparse training (DST) to all linear layers, excluding the embedding and final generator layer. Detailed hyperparameter settings for each experiment are provided in Tables 45, and 6. We also conduct a series of ablation and sensitivity tests on the components proposed in this article for CHTs and CHTss in Appendix H.
@@ 4.2 MLP for image classification
Table 2: Performance comparison on machine translation tasks of Multi30k, IWSLT, and WMT with varying final sparsity levels. The scores indicate BLEU scores, which is the higher the better. CHTs (GMP) represents CHTs with GMP’s density decay strategy. Bold values denote the best performance among fixed sparsity DST methods or density decay DST methods. The performances that surpass the fully connected model are marked with “*". The density decay of GMP, GraNet, and CHTss starts with a sparsity of 50%. The scores are averaged over three seeds ± their standard error.
@@
CHTs 28.94 ± 0.57 29.81 ± 0.37 21.15 ± 0.10 21.92 ± 0.17 20.94 ± 0.63 22.40 ± 0.06
MESTEM&S 28.89± 0.26 30.04 ± 0.52 19.56 ± 0.10 21.05 ± 0.21 20.70 ± 0.07 22.22 ± 0.10
GMP 30.51 ± 0.82 30.49 ± 0.40 22.76 ± 0.82 22.82 ± 0.53 22.47 ± 0.10 23.37 ± 0.08
GraNet 31.31 ± 0.31 31.62 ± 0.48* 22.53 ± 0.12 22.43 ± 0.09 22.51 ± 0.21 23.46 ± 0.09
CHTss 32.03 ± 0.29* 32.86 ± 0.16* 24.51 ± 0.02* 24.31 ± 0.04 23.73 ± 0.43 24.61 ± 0.14
@@
In the MLP evaluation, we aim to assess the fundamental capacity of DST methods to train the fully connected module, which is common across many ANNs. The sparse topological initialization of CHT and CHTs is CSTI zhang2023ultra since the input bipartite layer can directly receive information from the input pixels. Table 7 displays the performance of DST methods compared to their fully connected counterparts across three basic datasets of MNIST, Fashion MNIST, and EMNIST. The DST methods are tested at 99% sparsity. As shown in Table 7, both of the two regrowth methods of CHTs outperform the other fixed sparsity DST methods. Notably, the path-based CH3-L3p outperforms the fully connected one in all the datasets. The node-based CH2-L3n also achieves comparable performance on these basic datasets. However, considering the running time of CH3-L3p is unacceptable, especially in large scale models, in the rest of the experiments of this article, we only use CH2-L3n as the representative method to regrow new links. Table 1 presents a comparison of fixed-sparsity dynamic sparse training (DST) methods against the fully connected (FC) baseline. Notably, CHTs outperforms all other DST methods and achieve an 11% improvement in accuracy over the fully connected model. In addition, we present the active neuron post-percolation rate (ANP) for each method in Table 7 and Table 1. It is evident that CHTs adaptively percolates the network more effectively while retaining performance.